𝑓(π‘₯) = π‘₯

3 βˆ’ 12π‘₯, π‘₯ ∈ [0, 4]


1
Expert's answer
2021-05-07T10:30:51-0400

Ans:- "f(x)=x^3-12x" , "\ud835\udc65 \u2208 [0, 4]"

"\Rightarrow" "f(x)=x(x^2-12)\\"

"\Rightarrow f(x)=x(x+2\sqrt{3})(x-2\sqrt3)"

"\Rightarrow f(x)=x(x+3.4641)(x-3.4641)"

We know that "x" lies in the interval between "\ud835\udc65 \u2208 [0, 4]".

Hence The values of "x" for which "f(x)" should be zero is "0, 3.4641" and

"f'(x)=3x^2-12"

"\Rightarrow" "f'(0)=-12<0" so at "x=0" the slope of the f(x) is increasing

"\Rightarrow f'(3.4641)=-5.5692<0" so at "x=3.4641" the slope of the "f(x)" is increasing